OpenStudy (anonymous):

Please help me..... =*( Line A has slope of -5/3 and passes through the point (-2,7). What is the x-intercept of line A?

OpenStudy (zale101):

First, use the given infiltration to construct the equation of line A, then set y equals to zero and solve for x. The x's that you solved for are the x-intercepts.

OpenStudy (anonymous):



OpenStudy (zale101):

You can use the point slope formula and plug all the given info in. Do you remember the point slope formula?

OpenStudy (anonymous):

I have a problem understanding where to put the numbers... In what cases do you replace it for Y and in what cases do u replace it for b?

OpenStudy (anonymous):

y=mx+b

OpenStudy (zale101):

\(\Large y-y_1=m(x-x_1)\) Where x1,y1 is the given point and m is the slope

OpenStudy (zale101):

That is the point slope formula ^^

OpenStudy (anonymous):

is there any way to solve the problem without the point slope formula? because the solution doesnt use it

OpenStudy (zale101):

Or you can use y=mx+b Plug x,y by the given point and m for the slope Then solve for b After solving for b, you then go back and write y=mx+b again and plug in the y-intercept (b) and slope. Then, you have the equation solved.

OpenStudy (anonymous):

yea thats what i have a problem understanding... in what case do you use it for b and in what case do you use it for Y???????????

OpenStudy (zale101):

All the matters is to construct the equation of line A so you can set y to zero and solve for the x-intercept. There's two ways you can do it.

OpenStudy (anonymous):

i really need to understand that for future problems

OpenStudy (zale101):

Yes, i understand.